OpenStudy (anonymous):

What is the midpoint of the line segment with endpoints (3.2, 2.5) and (1.6, -4.5)?

OpenStudy (anonymous):

midpoint formula is \[\frac{ x1+x2 }{ 2 } ; \frac{ y2+y1 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 3.2+1.6 }{ 2};\frac{ 2.5+(-4.5) }{ 2 }\]

OpenStudy (adilalvi):

2,4,-1,5